Company XYZ uses a standard costing to evaluate operational performance against budget.

The following budget and actual information is presented to you as the accountant:

Budget

Production and sales:                    50 000 units

Material usage per unit:              7kg

Price per kg:                                     R10

Labour minutes per unit:              15 min

Total direct labour cost:                R2 000 000

Variable costs:                                R1 500 000

Actual

Production and sales:                    47 500 units

Material used:                                356 250kg

Cost of material:                             R3 241 875

Labour minutes:                             703 000 min

Total direct labour cost:                R1 900 000

Variable costs:                                R1 400 000

REQUIRED:

Identify and apply an appropriate variance analysis to evaluate direct material, direct labour and variable overheads (where applicable, express answers to one decimal place). Please provide possible reasons for the discrepancies.                         
Answer #1
Budget Standard quantity (units) 50000.0
Material usage per unit (kg) 7.0
Standard price per kg 10.0
Standard quantity (in kg) 350000.0
Actual Production and sales (units) 47500.0
Material used: 356250.0
Cost of material: 3241875.0
Material Usage Variance: Computation
Standard quantity (in kg) *Standard price per kg =(350000*10) 3500000.0
Actual quantity used (in kg)*Standard price per kg =(356250*10) 3562500.0
Variance -62500.0
Reason Variance is adverse as the actual quantity consumed is more than the budgeted quantity
Material Price Variance: Computation
Actual quantity used (in kg)*Standard price per kg =(356250*10) 3562500.0
Actual Cost of material: 3241875.0
Variance 320625.0
Reason Variance is favourable since the actual price is less than the budgeted price for actual quantity.
Budget Total direct labour cost 2000000.0
Actual Total direct labour cost 1900000.0
Labour Cost Variance 100000.0
Reason Variance is favourable since the actual cost is less than the budgeted cost of labour.
Budget Variable costs: 1500000.0
Actual Variable costs: 1400000.0
Variable Cost Variance 100000.0
Reason Variance is favourable since the actual cost is less than the budgeted cost for variable overhead.